40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5055     Entry time: Thu Jul 28 16:26:42 2011
Author: steve 
Type: Update 
Category: SUS 
Subject: ITMX, ITMY & ETMY OSEM gains adjusted 

 

OSEM damping gains were adjusted by observing  real time dataviewer to get Q of 5

OSEMs were kicked up one by one with 200 counts  ~1sec. The error signal was optimized to get 1/2 of exitation amplitude at the 5th sinusoid wave.

C1: SUS-ITMX_SUSPOS_N1 gain   111      ->        65,             PIT        7.2      ->    8,           YAW        12      ->     6,          SIDE     280

                  ITMY                                      277       ->     120,                        19.2     ->     7,                             24     ->     19,                        420    ->      470

                  ETMY                                       10        ->       32,                          20      ->     3,                             20     ->     10,                          50

                  ETMX                                       22        ->      25,                             3,                                                3,                                              -170  

ETMX having problems:  1, YAW can not be excited

                                             2, SIDE has no error signal in dataviewer. Sensing voltage on MEDM screen 0.142V

 

                      

 

ELOG V3.1.3-